OpenStudy (anonymous):

in laplace transform unit step function is 1/s it is more complicated than 1. please help me

OpenStudy (across):

This makes no sense.

OpenStudy (across):

\[\mathcal{U}(t-a)\stackrel{\mathcal{L}}{\longrightarrow}\frac{e^{-as}}{s}.\]Is this what you're referring to?

OpenStudy (anonymous):

a function is attached with e^-st and integrate zero to infinity , we can use any letter insted of s what is the significance of s

OpenStudy (experimentx):

the Laplace transform of unit step \( \theta(t-0) \) is 1so ??

OpenStudy (lgbasallote):

s is just a variable.....

OpenStudy (experimentx):

*1/s

OpenStudy (experimentx):

try this \[ \int_0^{\infty} \theta(t) e^{-st} dt\] this is same as Laplace transform of unit value because laplace transform naturally chops off any information that is before the time t=0.

OpenStudy (experimentx):

the step function \( \theta(t-0)\) naturally chops off into that is at <t=0 ... so it would be same.